I need to rewrite the equation so that y is a function of x but I do not know how to do this with a fraction.
1/4y+3=-5X
This is the answer I have but I am pretty sure it is not right. Could you please explain?
y = -5x/1/4-3/1/4
I think maybe I should multiply the fraction.





Answer by Theo(2967)

1/4y+3=-5x

Subtract 3 from both sides to get:

(1/4)y = -5x - 3

Multiply both sides by 4 to get:

y = 4*(-5x-3)

This becomes:

y = -20x - 12

Your answer was:

y = -5x/1/4-3/1/4

which I think was meant to be:

y = -5x/(1/4) - 3/(1/4)

1/(1/4) is the same as 1*4 so your answer is the same as mine.

Your answer becomes:

y = -5x / (1/4) - 3 / (1/4) which becomes:

y = -5x * 4 - 3 * 4 which becomes:

y = -20x - 12

I usually multiply to clear the fraction.

Makes solving the problem a little easier for me.

Either way gets the right answer.